OpenStudy (anonymous):

Use properties to evaluate the expression. (2 • 9) •4 A.72......B.54.........C.38...........D.15

TheSmartOne (thesmartone):

Do you know what \(2\times9=?\)

TheSmartOne (thesmartone):

@Freedom101

TheSmartOne (thesmartone):

\[(a \times b)\times c=a \times (b \times c)=a \times b \times c\]

TheSmartOne (thesmartone):

So in your case what does \(2 \times 9 \times 4=?\)